# The Fundamental Parameters of Astrophysical Plasma Turbulence and its Dissipation: Nonrelativistic Limit

### v1

@@ 5.1 Proposed Damping Mechanisms for Kinetic Turbulence
where $\Omega_{s}$ is the cyclotron frequency for species $s$ and $|n|>1$ indicates the harmonics of the cyclotron frequency. The $n=0$ resonance is often denoted the Landau resonance and includes two damping mechanisms: (i) Landau damping is mediated by the parallel component of the electric field $E_{\parallel}$ doing work on the charged particles (Landau, 1946; Leamon et al., 1998b, a; Quataert, 1998; Leamon et al., 1999; Quataert & Gruzinov, 1999; Leamon et al., 2000; Howes et al., 2008a; Schekochihin et al., 2009; TenBarge & Howes, 2013; Howes, 2015; Li et al., 2016; Chen et al., 2019; Afshari et al., 2021); and (ii) transit-time damping, also known as Barnes damping, is mediated by the magnetic mirror force associated with magnetic field magnitude fluctuations doing work on the magnetic moment of the charged particle gyromotion (Barnes, 1966; Quataert, 1998; Quataert & Gruzinov, 1999; Howes et al., 2008a; Huang et al., 2024). Both of these Landau-resonant collisionless damping mechanisms energize particles in the degree of freedom parallel to the local magnetic field. The $n\neq 0$ resonances yield cyclotron damping of the turbulent flucutations by $E_{\perp}$, energizing particles in the two perpendicular degrees of freedom (Coleman, 1968; Denskat et al., 1983; Isenberg & Hollweg, 1983; Goldstein et al., 1994; Leamon et al., 1998a; Gary, 1999; Isenberg et al., 2001; Hollweg & Isenberg, 2002; Isenberg & Vasquez, 2019; Afshari et al., 2023).
@@
Non-resonant damping mechanisms can remove energy from turbulent fluctuations through wave-particle interactions that do not require satisfying a particular resonance condition. Stochastic heating can arise when sufficiently large amplitude turbulent electromagnetic fluctuations can effectively scatter the otherwise organized gyromotion of particles, leading to a heating of the particles in the perpendicular degrees of freedom (Johnson & Cheng, 2001; Chen et al., 2001; White et al., 2002; Voitenko & Goossens, 2004; Bourouaine et al., 2008; Chandran et al., 2010; Chandran, 2010; Chandran et al., 2011; Bourouaine & Chandran, 2013; Chandran et al., 2013; Klein & Chandran, 2016; Arzamasskiy et al., 2019; Martinović et al., 2020; Cerri et al., 2021). Magnetic pumping energizes particles through a combination of two effects: (i) particles undergo the reversible, double adiabatic evolution of their perpendicular and parallel velocities in a magnetic field with a time varying magnitude; and (ii) those particles are scattered by collisions during this evolution, introducing irreversibility and leading to a small net transfer of energy to the particles over a full oscillation cycle (Spitzer & Witten, 1951; Berger et al., 1958; Lichko et al., 2017; Lichko & Egedal, 2020; Montag & Howes, 2022). One final proposed non-resonant damping mechanism is the recently identified kinetic “viscous” heating mediated by the effective collisionality associated with temperature anisotropy instabilities driven by large-scale compressible fluctuations in high beta plasmas (Arzamasskiy et al., 2023).
@@ 5.2 A Competition between Turbulent Cascade Rates and Damping Rates
One of the most challenging aspects of modeling the damping of plasma turbulence is that the net energy transfer to each species arises from the competition between the turbulent cascade rate and the kinetic damping rates as a function of scale, as depicted by the arrows in Fig. 1(a). Nonlinear interactions among turbulent fluctuations—denoted Alfvén wave collisions (Howes & Nielson, 2013) in the case of Alfvénic turbulence—mediate the transfer of energy from large to small scales that is dominantly local in scale space (Howes et al., 2011; Told et al., 2015), where energy is transferred locally from a wavenumber $k$ to a wavenumber $2k$, then on to $4k$, and so on. This is indicated by the black curved arrows in the figure. Because the physics of the cascade is dominantly local in nature (Howes et al., 2011; Told et al., 2015), scaling theory implies that the energy transfer rate at a given wavenumber $k$ is determined by the conditions at that scale. Note that, when analyzed in Fourier space, the energy transfer to small scales in the turbulent cascade is the sum of the nonlinear interactions among many triads of plane-wave modes, with a wide spread of energy transfer rates about zero; the net energy transfer given by this sum at a given scale is much smaller in amplitude than the spread of individual three-wave interaction rates, and the sum typically indicates a forward cascade of energy to small scales (Coburn et al., 2014, 2015). The turbulent cascade rate $\varepsilon$ used here specifically refers to this net energy rate obtained by integrating over all contributing nonlinear interactions involving that scale.
@@
Figure 4: Diagram of the balance of the energy transfer rates in a steady-state turbulent cascade at a local wavenumber $k_{*}$, where the following terms sum to zero: (i) the turbulent nonlinear energy transfer rate from lower wavenumbers $\varepsilon(k^{-})^{(k^{-}<k_{*})}$; (ii) the turbulent nonlinear energy transfer rate to higher wavenumbers $\varepsilon(k_{*})^{(k_{*}<k^{+})}$; and (iii) the local (in scale) kinetic damping rates of the turbulent energy, due to the sum of ion energization rates $Q_{i}(k_{*})$ (blue) and electron energization rates $Q_{e}(k_{*})$ (magenta).
@@
Competing with that turbulent cascade rate at each scale is the damping rate due to collisionless interactions between the electromagnetic fields and the plasma particles, as shown by the blue arrows for the transfer of energy to ions and the magenta arrows for the transfer of energy to electrons in Fig. 1(a). The net damping rate of the entire turbulent cascade, resulting from the sum of the energization rates for each particle species $s$, can be computed by integration over all scales of the turbulent cascade. In a steady-state, at each wavenumber $k_{*}$, the sum of three terms must balance to zero, as illustrated in Fig. 4: (i) the energy transfer from lower wavenumbers $k^{-}<k_{*}$ (larger scales), denoted by $\varepsilon(k^{-})^{(k^{-}<k_{*})}$; (ii) the energy transfer to higher wavenumbers $k^{+}>k_{*}$ (smaller scales), denoted by $\varepsilon(k_{*})^{(k_{*}<k^{+})}$; and (iii) the kinetic damping of the fluctations at wavenumber $k_{*}$, given by $Q(k_{*})$, which is the sum of the local (in scale) ion damping $Q_{i}(k_{*})$ (blue) and electron damping $Q_{e}(k_{*})$ (magenta). Complicating the determination of the partitioning of damped turbulent energy between ion and electrons $Q_{i}/Q_{e}$ is the fact that the kinetic damping rates due to ions and electrons often overlap as a function of scale (Howes et al., 2011, 2011; Told et al., 2015).
@@
Determining this balance of cascade and damping rates as a function of scale, and ultimately using that to determine the partitioning of turbulent energy among species and degrees of freedom—parameterized by $Q_{i}/Q_{e}$, $Q_{\perp,i}/Q_{i}$, and $Q_{\perp,e}/Q_{e}$—has been attempted in empirical turbulent cascade models (Pao, 1965; Howes et al., 2008a; Podesta et al., 2010; Howes et al., 2011; Zhao et al., 2013; Schreiner & Saur, 2017). An example of the modeling of the turbulent energy cascade and its kinetic dissipation is presented in Fig. 5 from the results of the weakened cascade model by Howes et al. (2011). In addition to balancing the physics of the local cascade rate $\varepsilon(k_{\perp})$ with the kinetic damping rates due to ions $Q_{i}(k_{\perp})$ and electrons $Q_{e}(k_{\perp})$ via the Landau resonance, this cascade model also includes the contributions to the local cascade rate from nonlocal interactions by large-scale shearing and small-scale diffusion. The model results shown here in Fig. 5 are from the same calculation as presented in Fig. 8 of Howes et al. (2011), where we reproduce (a) the one-dimensional perpendicular magnetic energy spectrum $E_{B_{\perp}}(k_{\perp})$ and (b) the anisotropic cascade through wave vector space characterized by the characteristic parallel wavenumber as a function of the perpendicular wavenumber, $k_{\parallel}(k_{\perp})$. This model for a plasma with $\beta_{i}=1$ and isotropic equilibrium Maxwellian velocity distributions with $T_{i}/T_{e}=9$ is driven weakly with $\chi(k_{\perp 0})=0.1$ at $k_{\perp 0}\rho_{i}=k_{\parallel}\rho_{i}=10^{-5}$, taking the model Kolmogorov constants to be $C_{1}=1.4$ and $C_{2}=1.0$.
@@
The energy transfer rates $Q_{i}(k_{\perp})$ and $Q_{e}(k_{\perp})$ that remove energy from the turbulent fluctuations and transfer it to the plasma particles in turbulent cascade models require as input the kinetic damping rates due to ions $\gamma_{i}(k_{\perp})$ and electrons $\gamma_{e}(k_{\perp})$. In the case of the weakened cascade model shown in Fig. 5, the damping rates used are collisionless damping rates via the Landau ($n=0$) resonance for the Alfvén wave mode (and its extension as a kinetic Alfvén wave at scales $k_{\perp}\rho_{i}\gtrsim 1$) from the Vlasov-Maxwell linear dispersion relation (Howes et al., 2011). Previous studies have called into question whether resonant collisionless wave-particle interactions, such as Landau damping, can effectively damp turbulent fluctuations in the presence of the strong nonlinear interactions that mediate the turbulent cascade (Plunk, 2013; Schekochihin et al., 2016). Using the field-particle correlation technique (Klein & Howes, 2016; Howes et al., 2017; Klein et al., 2017), analysis of kinetic simulations (Klein et al., 2017; Howes et al., 2018; Klein et al., 2020) and spacecraft observations (Chen et al., 2019; Afshari et al., 2021, 2023) of plasma turbulence find clear velocity-space signatures of Landau damping. These findings indicate indeed that resonant wave-particle interactions do play a role, and possibly a dominant role (Afshari et al., 2021), in the damping of space and astrophysical plasma turbulence, settling the question of whether Landau damping can effectively damp strong plasma turbulence.
In quantifying how a particular damping mechanism competes with the nonlinear turbulent cascade, the key dimensionless quantity to evaluate is the ratio of linear damping rate due to nonlinear cascade rate $\gamma/\omega_{nl}$ (Howes et al., 2008a, 2011). Adopting the conjecture of critical balance between linear and nonlinear timescales in strong plasma turbulence $\omega\sim\omega_{nl}$ (Goldreich & Sridhar, 1995; Howes et al., 2008a, 2011; Mallet et al., 2015), the key dimensionless measure of the importance of that particular mechanism is the ratio $\gamma/\omega$, which can be estimated using the linear theory for each of the different damping mechanisms.
@@ 5.3 The Relevance of Linear Physics to Strong Plasma Turbulence
The concept of critical balance in plasma turbulence (Goldreich & Sridhar, 1995; Mallet et al., 2015)—which suggests that the nonlinear timescales of the turbulent cascade remain in approximate balance with the timescales of the linear wave physics, $\omega_{nl}\sim\omega$—implies that the physics of the linear plasma response remains valid in strong plasma turbulence. But, perhaps in analogy with the case of incompressible hydrodynamic turbulence in which no linear response exists (Howes, 2015), the validity of linear physics has been called into question in the presence of strong plasma turbulence (Matthaeus et al., 2014). Although there exist numerous counterexamples in which linear physics properties have been shown to be relevant to strong plasma turbulence (Maron & Goldreich, 2001; Cho & Lazarian, 2003; Alexandrova et al., 2008; Howes et al., 2008b, a; Svidzinski et al., 2009; Sahraoui et al., 2010; Hunana et al., 2011; Howes et al., 2011; TenBarge et al., 2012; Chen et al., 2013), the question has persisted. A more recent careful study of kinetic numerical simulations and spacecraft observations has shown clearly that both small-amplitude and large-amplitude fluctuations, including variations interpreted to be coherent structures, preserve the linear properties of Alfvénic fluctuations (Grošelj et al., 2019), hopefully providing sufficient evidence to establish that relevance of linear physics to strong plasma turbulence.
@@
In the limit of isotropic equilibrium velocity distributions for the ions and electrons—$A_{i}=1$, $A_{e}=1$, and $\tau=T_{i}/T_{e}$—this set of dependencies is reduced to
@@
For both the isotropic and anisotropic equilibirum temperatures, in the non-relativistic limit $v_{ti}/c\ll 1$ typical of many space and astrophysical plasmas of interest, the linear Vlasov-Maxwell dispersion relation is practically independent of $v_{ti}/c$, so this parameter may effectively be dropped for the case of non-relativistic plasma turbulence.
In the limit of a sufficiently large inertial range of turbulence, the MHD turbulence scaling parameters given by (1)–(4) predict that turbulent flucuations at the small-scale end of the inertial range, where the kinetic damping mechanisms are believed to become significant, will be anisotropic with $k_{\parallel}\ll k_{\perp}$ and small amplitude relative to the equilibrium magnetic field $|\delta\mathbf{B}|\ll B_{0}$ (Howes et al., 2006, 2008a; Schekochihin et al., 2009). In this limit, the turbulent frequencies of the incompressible Alfvénic fluctuations directly observed to dominate turbulence in space plasmas (Belcher & Davis, 1971; Tu & Marsch, 1995; Alexandrova et al., 2008; Howes et al., 2012a; Bruno & Carbone, 2013) have frequencies much smaller than the ion cyclotron frequency, $\omega/\Omega_{i}=k_{\parallel}d_{i}\ll 1$, where $d_{i}=v_{A}/\Omega_{i}=\rho_{i}/\sqrt{\beta_{i}}$ is the ion inertial length. In this limit, the turbulent fluctuations at the small-scale end of the inertial range satisfy the conditions for the applicabilty of gyrokinetic theory (Rutherford & Frieman, 1968; Taylor & Hastie, 1968; Catto, 1978; Antonsen & Lane, 1980; Catto et al., 1981; Frieman & Chen, 1982; Dubin et al., 1983; Hahm et al., 1988; Brizard, 1992; Howes et al., 2006; Brizard & Hahm, 2007; Schekochihin et al., 2009). Gyrokinetic theory can be rigorously derived from plasma kinetic theory under the limits of low frequency ($\omega/\Omega_{i}\ll 1$) and anisotropic $k_{\parallel}\ll k_{\perp}$ fluctuations in the non-relativisitc limit $v_{ti}/c\ll 1$. In the gyrokinetic approximation, the physics of the the cyclotron resonances and of the fast magnetosonic fluctuations—and their kinetic extension to whistler fluctuations (Howes et al., 2014)—are ordered out of the system. Gyrokinetics retains the physics of the $\mathbf{E}\times\mathbf{B}$ nonlinearity that mediates the turbulent cascade to small scales, full finite-Larmor-radius effects, and the collisionless damping associated with the $n=0$ Landau resonance. In this reduced limit, the complex eigenfrequencies of the linear gyrokinetic dispersion relation for isotropic equilibrium velocity distributions are linearly dependent on the parallel wavenumber, and have the significantly reduced dimensionless parameter dependency given by (Howes et al., 2006)
@@
Although the ion-to-electron mass ratio $\mu=m_{i}/m_{e}$ is not a parameter that varies in the physical world, many numerical studies of plasma turbulence choose to employ a reduced mass ratio for computational efficiency. Therefore, we choose to include the $\mu$ dependence of different proposed turbulent damping mechanisms in our calculations below to determine how the use of a reduced mass ratio may impact the partitioning of turbulent energy removed by different damping mechanisms. Note also that it is important to employ a sufficiently large mass ratio to achieve a separation between the ion and electron scales, with a previous study suggesting $m_{i}/m_{e}\geq 32$ is necessary to model the distinct ion and electron responses to the electromagnetic fluctuations due to this scale separation (Howes et al., 2018).

### v2

@@ 5.1 Proposed Damping Mechanisms for Kinetic Turbulence
where $\Omega_{s}$ is the cyclotron frequency for species $s$ and $|n|>1$ indicates the harmonics of the cyclotron frequency. The $n=0$ resonance is often denoted the Landau resonance and includes two damping mechanisms: (i) Landau damping is mediated by the parallel component of the electric field $E_{\parallel}$ doing work on the charged particles (Landau, 1946; Leamon et al., 1998b, a; Quataert, 1998; Leamon et al., 1999; Quataert & Gruzinov, 1999; Leamon et al., 2000; Howes et al., 2008a; Schekochihin et al., 2009; TenBarge & Howes, 2013; Howes, 2015; Li et al., 2016; Chen et al., 2019; Afshari et al., 2021); and (ii) transit-time damping, also known as Barnes damping, is mediated by the magnetic mirror force associated with magnetic field magnitude fluctuations doing work on the magnetic moment of the charged particle gyromotion (Barnes, 1966; Quataert, 1998; Quataert & Gruzinov, 1999; Howes et al., 2008a; Huang et al., 2024; Howes et al., 2024). Both of these Landau-resonant collisionless damping mechanisms energize particles in the degree of freedom parallel to the local magnetic field. The $n\neq 0$ resonances yield cyclotron damping of the turbulent fluctuations by $E_{\perp}$, energizing particles in the two perpendicular degrees of freedom (Coleman, 1968; Denskat et al., 1983; Isenberg & Hollweg, 1983; Goldstein et al., 1994; Leamon et al., 1998a; Gary, 1999; Isenberg et al., 2001; Hollweg & Isenberg, 2002; Isenberg & Vasquez, 2019; Afshari et al., 2023).
@@
Non-resonant damping mechanisms can remove energy from turbulent fluctuations through wave-particle interactions that do not require satisfying a particular resonance condition. Stochastic heating can arise when sufficiently large amplitude turbulent electromagnetic fluctuations can effectively scatter the otherwise organized gyromotion of particles, leading to a heating of the particles in the perpendicular degrees of freedom (Johnson & Cheng, 2001; Chen et al., 2001; White et al., 2002; Voitenko & Goossens, 2004; Bourouaine et al., 2008; Chandran et al., 2010; Chandran, 2010; Chandran et al., 2011; Bourouaine & Chandran, 2013; Chandran et al., 2013; Klein & Chandran, 2016; Arzamasskiy et al., 2019; Martinović et al., 2020; Cerri et al., 2021). Magnetic pumping energizes particles through a combination of two effects: (i) particles undergo the reversible, double adiabatic evolution of their perpendicular and parallel velocities in a magnetic field with a time varying magnitude; and (ii) those particles are scattered by collisions during this evolution, introducing irreversibility and leading to a small net transfer of energy to the particles over a full oscillation cycle (Spitzer & Witten, 1951; Berger et al., 1958; Lichko et al., 2017; Lichko & Egedal, 2020; Montag & Howes, 2022). One final proposed non-resonant damping mechanism is the recently identified kinetic “viscous” heating mediated by the effective collisionality associated with temperature anisotropy instabilities driven by large-scale fluctuations of the magnetic field magnitude in high beta plasmas (Arzamasskiy et al., 2023).
@@ 5.2 A Competition between Turbulent Cascade Rates and Damping Rates
One of the most challenging aspects of modeling the damping of plasma turbulence is that the net energy transfer to each species arises from the competition between the turbulent cascade rate and the kinetic damping rates as a function of scale, as depicted by the arrows in Fig. 1(a). Nonlinear interactions among turbulent fluctuations—denoted Alfvén wave collisions (Howes & Nielson, 2013) in the case of Alfvénic turbulence—mediate the transfer of energy from large to small scales that is dominantly local in scale space (Howes et al., 2011; Told et al., 2015), where energy is transferred locally from a wavenumber $k$ to a wavenumber $2k$, then on to $4k$, and so on. This is indicated by the black curved arrows in the figure. Because the physics of the cascade is dominantly local in nature (Howes et al., 2011; Told et al., 2015), scaling theory implies that the energy transfer rate at a given wavenumber $k$ is determined solely by the conditions at that scale. Note that, when analyzed in Fourier space, the energy transfer to small scales in the turbulent cascade is the sum of the nonlinear interactions among many triads of plane-wave modes, with a wide spread of energy transfer rates about zero; the net energy transfer given by this sum at a given scale is much smaller in amplitude than the spread of all of the individual three-wave interaction rates, and the sum typically indicates a forward cascade of energy to small scales for turbulence in three spatial dimensions (Coburn et al., 2014, 2015). The turbulent cascade rate $\varepsilon$ used here specifically refers to this net energy transfer rate obtained by integrating over all contributing nonlinear interactions involving that scale [^7] Because the net energy transfer rate at a given scale fundamentally involves the integration over all possible nonlinear triads, it can in principle include a non-negligible contribution from nonlocal interactions, such as those incorporated in the weakened cascade model of the turbulent cascade (Howes et al., 2011)..
@@
Figure 4: Diagram of the balance of the energy transfer rates in a steady-state turbulent cascade at a local wavenumber $k_{*}$, where the following terms sum to zero: (i) the turbulent nonlinear energy transfer rate from lower wavenumbers $\varepsilon(k^{-})^{(k^{-}<k_{*})}$; (ii) the turbulent nonlinear energy transfer rate to higher wavenumbers $\varepsilon(k_{*})^{(k_{*}<k^{+})}$; and (iii) the local (in scale) kinetic damping rates of the turbulent energy, due to the sum of ion energization rate $Q_{i}(k_{*})$ (blue) and electron energization rate $Q_{e}(k_{*})$ (magenta).
@@
Competing with that turbulent cascade rate at each scale is the damping rate due to collisionless interactions between the electromagnetic fields and the plasma particles, as shown by the blue arrows for the transfer of energy to ions and the magenta arrows for the transfer of energy to electrons in Fig. 1(a). The net damping rate of the entire turbulent cascade, resulting from the sum of the energization rates for each particle species $s$, can be computed by integration over all scales of the turbulent cascade. In a steady-state, at each wavenumber $k_{*}$, the sum of three terms must balance to zero, as illustrated in Fig. 4: (i) the energy transfer from lower wavenumbers $k^{-}<k_{*}$ (larger scales), denoted by $\varepsilon(k^{-})^{(k^{-}<k_{*})}$; (ii) the energy transfer to higher wavenumbers $k^{+}>k_{*}$ (smaller scales), denoted by $\varepsilon(k_{*})^{(k_{*}<k^{+})}$; and (iii) the kinetic damping of the fluctuations at wavenumber $k_{*}$, given by $Q(k_{*})$, which is the sum of the local (in scale) ion energization rate $Q_{i}(k_{*})$ (blue) and electron energization rate $Q_{e}(k_{*})$ (magenta). Complicating the determination of the partitioning of damped turbulent energy between ions and electrons $Q_{i}/Q_{e}$ is the fact that the kinetic damping rates due to ions and electrons may overlap as a function of scale (Howes et al., 2011, 2011; Told et al., 2015).
@@
Determining this balance of cascade and damping rates as a function of scale, and ultimately using the terms in that balance to determine the partitioning of turbulent energy among species and degrees of freedom—parameterized by $Q_{i}/Q_{e}$, $Q_{\perp,i}/Q_{i}$, and $Q_{\perp,e}/Q_{e}$—has been attempted in empirical turbulent cascade models (Pao, 1965; Howes et al., 2008a; Podesta et al., 2010; Howes et al., 2011; Zhao et al., 2013; Schreiner & Saur, 2017). An example of the modeling of the turbulent energy cascade and its kinetic dissipation is presented in Fig. 5 from the results of the weakened cascade model by Howes et al. (2011). In addition to balancing the physics of the local cascade rate $\varepsilon(k_{\perp})$ with the kinetic damping rates due to ions $Q_{i}(k_{\perp})$ and electrons $Q_{e}(k_{\perp})$ via the Landau resonance, this cascade model also includes the contributions to the local cascade rate from nonlocal interactions by large-scale shearing and small-scale diffusion. The model results shown here in Fig. 5 are from the same calculation as presented in Fig. 8 of Howes et al. (2011), where we reproduce (a) the one-dimensional perpendicular magnetic energy spectrum $E_{B_{\perp}}(k_{\perp})$ and (b) the anisotropic cascade through wave vector space characterized by the characteristic parallel wavenumber as a function of the perpendicular wavenumber, $k_{\parallel}(k_{\perp})$. This model for a plasma with $\beta_{i}=1$ and isotropic equilibrium Maxwellian velocity distributions with $T_{i}/T_{e}=9$ is driven weakly with $\chi(k_{\perp 0})=0.1$ at $k_{\perp 0}\rho_{i}=k_{\parallel}\rho_{i}=10^{-5}$, taking the model Kolmogorov constants to be $C_{1}=1.4$ and $C_{2}=1.0$. Note that an important avenue of research is to use kinetic numerical simulations or spacecraft observations to constrain the values of these Kolmogorov constants, as recently done by Shankarappa et al. (2023) using observations in the inner heliosphere by the Parker Solar Probe (Fox et al., 2016).
@@
The energy density transfer rates $Q_{i}(k_{\perp})$ and $Q_{e}(k_{\perp})$ that remove energy from the turbulent fluctuations and transfer it to the plasma particles in turbulent cascade models require as input the kinetic damping rates due to ions $\gamma_{i}(k_{\perp})$ and electrons $\gamma_{e}(k_{\perp})$. In the case of the weakened cascade model shown in Fig. 5 (Howes et al., 2011), the damping rates used are collisionless damping rates via the Landau ($n=0$) resonance for the Alfvén wave mode (and its extension as a kinetic Alfvén wave at scales $k_{\perp}\rho_{i}\gtrsim 1$) from the Vlasov-Maxwell linear dispersion relation (Stix, 1992). Previous studies have called into question whether resonant collisionless wave-particle interactions, such as Landau damping, can effectively damp turbulent fluctuations in the presence of the strong nonlinear interactions that mediate the turbulent cascade (Plunk, 2013; Schekochihin et al., 2016). Using the field-particle correlation technique (Klein & Howes, 2016; Howes et al., 2017; Klein et al., 2017), several recent analyses of kinetic simulations (Klein et al., 2017; Howes et al., 2018; Klein et al., 2020) and spacecraft observations (Chen et al., 2019; Afshari et al., 2021) of plasma turbulence find clear velocity-space signatures of Landau damping. These findings indicate indeed that resonant wave-particle interactions do play a role, and possibly a dominant role (Afshari et al., 2021), in the damping of space and astrophysical plasma turbulence, settling the question of whether Landau damping can effectively damp strong plasma turbulence [^8] Note that recent work found that an anti-phase-mixing process can inhibit the process of the linear phase mixing that transfers nonthermal internal energy to small parallel velocity-space scales, thereby shutting down the effect of Landau damping of compressible turbulent fluctuations within the inertial range under sufficiently collisionless conditions (Schekochihin et al., 2016; Parker et al., 2016; Meyrand et al., 2019)..
In quantifying how a particular damping mechanism competes with the nonlinear turbulent cascade, the key dimensionless quantity to evaluate is the ratio of linear damping rate due to nonlinear cascade rate $\gamma/\omega_{nl}$ (Howes et al., 2008a, 2011). Adopting the conjecture of critical balance between linear and nonlinear timescales in strong plasma turbulence $\omega\sim\omega_{nl}$ (Goldreich & Sridhar, 1995; Howes et al., 2008a, 2011; Mallet et al., 2015), the key dimensionless measure of the importance of any particular mechanism is the ratio $\gamma/\omega$, which can be estimated using the linear theory for each of the different damping mechanisms.
@@ 5.3 The Relevance of Linear Physics to Strong Plasma Turbulence
The concept of critical balance in plasma turbulence (Goldreich & Sridhar, 1995; Howes et al., 2008a, 2011; Mallet et al., 2015)—which suggests that the nonlinear timescales of the turbulent cascade remain in approximate balance with the timescales of the linear wave physics, $\omega_{nl}\sim\omega$—implies that the physics of the linear plasma response remains valid even in strong plasma turbulence. But, perhaps in analogy with the case of incompressible hydrodynamic turbulence in which no linear response exists (Howes, 2015), the validity of linear physics has been called into question in the presence of strong plasma turbulence (Matthaeus et al., 2014). Although there exist numerous counterexamples in which linear physics properties have been shown to be relevant to strong plasma turbulence (Maron & Goldreich, 2001; Cho & Lazarian, 2003; Alexandrova et al., 2008; Howes et al., 2008b, a; Svidzinski et al., 2009; Sahraoui et al., 2010; Hunana et al., 2011; Howes et al., 2011; TenBarge et al., 2012; Chen et al., 2013), the question has persisted. A more recent careful study of kinetic numerical simulations and spacecraft observations has shown clearly that both small-amplitude and large-amplitude fluctuations, including variations interpreted to be coherent structures, preserve the linear properties of Alfvénic fluctuations (Grošelj et al., 2019), hopefully providing sufficient evidence to establish definitively the relevance of linear physics to strong plasma turbulence.
@@
In the limit of isotropic equilibrium velocity distributions for the ions and electrons—leading to the simplifications $A_{i}=1$, $A_{e}=1$, and $\tau_{\parallel}=\tau\equiv T_{i}/T_{e}$—this set of dependencies is reduced to
@@
For both the isotropic and anisotropic equilibrium temperatures, in the non-relativistic limit $v_{ti}/c\ll 1$ typical of many space and astrophysical plasmas of interest, the linear Vlasov-Maxwell dispersion relation is practically independent of $v_{ti}/c$, so this parameter may effectively be dropped for the case of non-relativistic plasma turbulence.
In the limit of a sufficiently large inertial range of turbulence, the MHD turbulence scaling parameters given by (1)–(4) predict that turbulent fluctuations at the small-scale end of the inertial range—where most of the kinetic damping mechanisms are believed to become significant—will be anisotropic with $k_{\parallel}\ll k_{\perp}$ and small amplitude relative to the equilibrium magnetic field $|\delta\mathbf{B}|\ll B_{0}$ (Howes et al., 2006, 2008a; Schekochihin et al., 2009). In this limit, the turbulent frequencies of the incompressible Alfvénic fluctuations directly observed to dominate turbulence in space plasmas (Belcher & Davis, 1971; Tu & Marsch, 1995; Alexandrova et al., 2008; Howes et al., 2012a; Bruno & Carbone, 2013) have frequencies much smaller than the ion cyclotron frequency, $\omega/\Omega_{i}=k_{\parallel}d_{i}\ll 1$, where $d_{i}=v_{A}/\Omega_{i}=\rho_{i}/\sqrt{\beta_{i}}$ is the ion inertial length. In this limit, the turbulent fluctuations at the small-scale end of the inertial range satisfy the conditions for the applicability of gyrokinetic theory (Rutherford & Frieman, 1968; Taylor & Hastie, 1968; Catto, 1978; Antonsen & Lane, 1980; Catto et al., 1981; Frieman & Chen, 1982; Dubin et al., 1983; Hahm et al., 1988; Brizard, 1992; Howes et al., 2006; Brizard & Hahm, 2007; Schekochihin et al., 2009). Gyrokinetic theory can be rigorously derived from plasma kinetic theory under the limits of low frequency ($\omega/\Omega_{i}\ll 1$) and anisotropic ($k_{\parallel}\ll k_{\perp}$) fluctuations in the non-relativistic limit ($v_{ti}/c\ll 1$). In the gyrokinetic approximation, the physics of the the cyclotron resonances and of the fast magnetosonic fluctuations—and their kinetic extension to whistler fluctuations (Howes et al., 2014)—are ordered out of the system. Gyrokinetics retains the physics of the $\mathbf{E}\times\mathbf{B}$ nonlinearity that mediates the turbulent cascade to small scales, full finite-Larmor-radius effects, and the collisionless damping associated with the $n=0$ Landau resonance. In this reduced limit, the complex eigenfrequencies of the linear gyrokinetic dispersion relation for isotropic equilibrium velocity distributions are linearly dependent on the parallel wavenumber, and have the significantly reduced dimensionless parameter dependency given by (Howes et al., 2006)
@@
Although the ion-to-electron mass ratio $\mu=m_{i}/m_{e}$ is not a parameter that varies in the physical world, many numerical studies of plasma turbulence choose to employ a reduced mass ratio for computational efficiency. Therefore, we choose to include the $\mu$ dependence of different proposed turbulent damping mechanisms in our calculations below to determine how the use of a reduced mass ratio may impact the partitioning of turbulent energy removed by different damping mechanisms. Note also that it is important to employ a sufficiently large mass ratio to achieve a separation between the ion and electron scales, with a previous study suggesting $m_{i}/m_{e}\geq 32$ is necessary to model with qualitative fidelity the distinct ion and electron responses to the electromagnetic fluctuations due to this scale separation (Howes et al., 2018).